OpenStudy (anonymous):

What is the answer to this? Find the approximate solution of the initial value problem presented below in terms of a Taylor polynomial of degree 4. y' + x^2-y^2 = 0, y(0) = 1

OpenStudy (kainui):

Solve for y'(x) and you'll see that y'(x)=-x^2+y(x)^2 We know from the problem: y(0)=1 y'(0)=0+1^2=1 I'll help you get started on the next one even, y''(x)=-2x+2y(x)y'(x) y''(0)=? y'''(0)=? You just need to approximate it to 4 and then put it into a taylor polynomial. Don't forget the whole point of a taylor polynomial is to approximate them so that they have the same first, second, ... nth derivative at x=0.

OpenStudy (kainui):

I hope that helped... Now that I reread the question I'm not sure if it will or not lol.